A proton moving at 7.00 x 106 m/s through a magnetic field of magnitude 2.6 T experiences magnetic force of magnitude 6.9x10-13 N. What is the angle between the proton's velocity and the field?

6.2°

76.3°

13.7°

44.9°
Answer #1

B = 2.6T, F = 6.9x16 on Given 8- v=7.0x108 m/s q = 1.6810-19 C The magnetic force sinos 0 = sin(0:234) 0 = 13.7 Ans a proton13.7°